OpenStudy (anonymous):

suppose that xe^yz, where x=t2, y=1-t and z 1+2t find dw/dt

OpenStudy (dumbcow):

please clarify... where is w?

OpenStudy (dumbcow):

im guessing you meant x = t^2

OpenStudy (anonymous):

ya is t^2 and w= xe^y/z

OpenStudy (dumbcow):

just sub everything in so that w is in terms of "t" then differentiate

OpenStudy (anonymous):

t^2 e^1-t/ 1+2t like this?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

2t e^1-t/ 1+2t (-1 +2) =2t e^1-t/ 1+2t (1) ?

OpenStudy (dumbcow):

hard to read in that format...can you use equation editor? you have to use both product and quotient rules

OpenStudy (dumbcow):

here is solution to check your work http://www.wolframalpha.com/input/?i=differentiate++%28t^2+e^%281-t%29%29+%2F+%281%2B2t%29

OpenStudy (anonymous):

\[t^2 e^{1-t/1+2t} => 2t e^e^{1-t/1+2t} +t^2 e^{1-t/1+2t} (1-t/1+2t) \]